OpenStudy (anonymous):

If (2000!)/(1000!) = k (1 x 3 x 5 x 7 ... 1997 x 1999). What is the value of k?

myininaya (myininaya):

\[\frac{2000!}{1000!}=\frac{2000 \cdots \cdot 1001 \cdot 1000!}{1000!}=2000 \cdots 1001\] so we have \[2000 \cdots 1001 =k(1 \cdot 3 \cdot 5 \cdot 7 \cdots 1997 \cdot 1999)\] \[k=\frac{2000 \cdots 1001}{1 \cdot 3 \cdots 1999}\]

OpenStudy (anonymous):

I had trouble with simplifying.

OpenStudy (anonymous):

k = (2 x 4 x 6 x 8 x... 1998 x 2000)/ (1000!) (2(1 x 2 x 3 x 4 x... 999 x 1000)/ (1000!) = (2(1000!)/(1000!) = 2

OpenStudy (anonymous):

k = 2

OpenStudy (anonymous):

Sorry, but that does not seem correct. The denominator is solely composed of odd numbers.

OpenStudy (anonymous):

You know the answer?

OpenStudy (anonymous):

No. That's why I'm asking ... but 2 cannot be correct.

OpenStudy (anonymous):

2000! / (1 x 3 x 5 x 7 ... 1997 x 1999) = 2 x 4 x 6 x 8 x ... 2000

OpenStudy (anonymous):

It's not 2000!. The numbers only go down to 1001.

OpenStudy (anonymous):

k should be 2^1000 i believe

OpenStudy (anonymous):

@joemath 2 x 4 x 6 x 8 x ... 2000 = 2(1000!)

OpenStudy (anonymous):

no, (2+4+6+...+2000)=2(1+2+...+1000), we are multiplying here.

OpenStudy (zarkon):

joe is correct

OpenStudy (anonymous):

Yeah sorry

OpenStudy (anonymous):

I thought it was plus

OpenStudy (anonymous):

I think Joe has the correct answer.

myininaya (myininaya):

sorry joe i have to go don't be mad

myininaya (myininaya):

always leave when he comes on

myininaya (myininaya):

lol

myininaya (myininaya):

bye zarkon and everyone else

OpenStudy (zarkon):

later

OpenStudy (anonymous):

I think k = 2^1000 Mathematica solution is attached.

OpenStudy (zarkon):

we already have a solution...Mathematica is not needed...the human mind triumphs again :)

OpenStudy (anonymous):

Sorry. I was working on it and didn't notice the posting.